OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. This literally feels like 8+2. 2(4) + 2 8 + 2 10?

jimthompson5910 (jim_thompson5910):

if they asked `what is f(x) when x = 4` then you'd be correct

OpenStudy (anonymous):

Guess I'm wrong. XD

jimthompson5910 (jim_thompson5910):

they want you to find the inverse first, then plug in x = 4 into the inverse

OpenStudy (anonymous):

How do I find it again?

jimthompson5910 (jim_thompson5910):

step 1) replace f(x) with y step 2) swap x and y step 3) solve for y

jimthompson5910 (jim_thompson5910):

f(x) = 2x + 2 y = 2x + 2 x = 2y + 2 solve for y

OpenStudy (anonymous):

So... Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 2x = 2 2(4) = 2 8 = 2 4?

jimthompson5910 (jim_thompson5910):

if you solve `x = 2y + 2` for y, what do you get?

OpenStudy (anonymous):

xD 2 + 2 = 4 I think I'm trying to make it more complicated than it really is.

OpenStudy (anonymous):

Is that right? It's 4?

jimthompson5910 (jim_thompson5910):

step 1) replace f(x) with y step 2) swap x and y step 3) solve for y f(x) = 2x + 2 y = 2x + 2 ... step 1 x = 2y + 2 ... step 2 x-2 = 2y + 2-2 ... subtract 2 from both sides x-2 = 2y 2y = x-2 2y/2 = (x-2)/2 ... divide both sides by 2 y = (x-2)/2 So the inverse function is \[\Large f^{-1}(x) = \frac{x-2}{2}\]

jimthompson5910 (jim_thompson5910):

Now you plug x = 4 into \[\Large f^{-1}(x) = \frac{x-2}{2}\]